Evaluate the following: 
 
sin 60° cos 30° + sin 30° cos 60°

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

sin 60° cos 30° + sin 30° cos 60°
First, find the values of the given trigonometric ratios
sin 30° = 1/2
cos 30° =√3/2
sin 60° = 3/2
cos 60°= 1/2
Now, substitute the values in the given problem
sin 60° cos 30° + sin 30° cos 60°
=√3/2 ×√3/2+ (1/2) ×(1/2 )
= 3/4+1/4
= 4/4
=1

Thanks